# Efficient generative adversarial networks using linear additive-attention Transformers

### v4

@@ I Introduction
Generative Adversarial Networks (GANs) [1] have been one of the preferred approaches to image generation over the past decade [2, 3]. A key benefit of GANs lies in their ability to learn a smooth and continuous latent space representation of the data. This latent space allows for intuitive manipulation of the generated samples, such as smoothly transitioning between different samples and performing arithmetic operations on the latent vectors. Although diffusion models [4] have recently outperformed GANs in different generation tasks, they require multiple forward passes to generate a single sample, which translates into large computational resources. In contrast, GANs can generate samples in a single forward pass without additional training procedures, such as progressive distillation [5]. Moreover, some results suggest that GANs can match the quality of diffusion models (DM) using smaller networks (e.g. ADM-IP [6] with 296 million parameters and VITGAN [7] with 38 million parameters achieve a similar performance on CelebA).
Training instability is a challenge that has limited the application and adoption of GANs. Standard techniques to tackle this issue propose different objective functions (e.g. Wasserstein distance [8]) and regularization techniques (e.g. spectral normalization [9]). Moreover, laborious engineering and sophisticated neural modules [10, 11] are required to achieve state-of-the-art image generation using Conv-based GANs, resulting in computationally demanding architectures (e.g. StyleGANs [10, 11]) regarding parameters and FLOPs. Since the inductive bias of convolutions has been shown to limit long-range dependency learning [12], different GAN architectures that incorporate Transformers [13] have been proposed. However, self-attention can make GAN training even more unstable [7], and its $O(N^{2})$ complexity results in high computational demands [7, 14].
@@
We exhaustively evaluate the proposed model with different resolutions of CIFAR-10, CelebA, FFHQ, and LSUN Bedroom. LadaGAN achieves a competitive FID score compared with state-of-the-art GANs, ADM-IP and CT on all these datasets, while requiring significantly fewer FLOPs and parameters. In order to validate the effectiveness of the proposed architecture, we evaluate various $O(N)$ attention mechanisms with gradient penalty and their compatibility with the inductive bias of convolutions. Our findings indicate that the LadaGAN architecture is gradient-stable and its attention mechanism is convolution-compatible, thus proving to be a highly effective architecture for image generation.
@@ IV Experiments
Attention Convolutions Residual-MLP Modulation FLOPs FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ Prec $\uparrow$ Rec $\uparrow$
@@ IV-C Ablation studies
Table I shows the FID scores, IS Precision, Recall, and number of FLOPs for all the evaluated attention mechanisms and configurations. As can be observed, the Ladaformer with a convolutional layer and without the residual connection achieves the best evaluation with a similar number of FLOPs. In general, adding a convolutional layer has a positive effect on Ladaformer, a negative effect for Linformer and practically no effect for Fastformer. This shows not only that the combination of Ladaformer and convolutions does not lead to training instabilities, but also that it can provide noticeable benefits for the quality of the generated images. This is because the locality of the convolutional layer might complement the long-range dependencies of the additive attention map that is propagated by the element-wise operation in equation 3. Note that, as opposed to Lada, Fastformer attention mechanism compresses the representation by computing a second additive attention map for the keys instead of propagating it through an element-wise operation, which seems to prevent the benefits of the convolutional layer.
@@ IV-D Lada discriminator
Table II shows the FIDs obtained with different configurations of ConvNet and Ladaformer discriminators on CIFAR-10, CelebA and LSUN Bedroom. In addition, the number of FLOPs for the discriminator and generator, as well as the sizes of the embeddings in the generator corresponding to the three Ladaformer blocks in Figure 4, are also presented. As can be observed, the Ladaformer discriminators consistently outperform convolutional discriminators. However, the difference is small for CIFAR-10 and CelebA and significantly larger for LSUN Bedroom where larger generators are employed. Interestingly, a higher learning rate is required for the convolutional discriminator to match the Ladaformer discriminator FID. We also observe that, in LSUN Bedroom, using LEE leads to the lowest FID, while in CIFAR-10 and CelebA LEE does not make any difference. We hypothesize that this behavior is due to the resolution of the dataset, since high-resolution images require larger models. Note that in the LSUN Bedroom experiment, the Ladaformer blocks of the generator increase their embedding dimension locally from $\{1024,256,64\}$ to $\{1024,512,256\}$, which in turn increases the number of FLOPs. deteriorates
@@ IV-F Comparison with state-of-the-art models
Resolution CT [33] ADM-IP [6] StyleGAN2 [11] VITGAN [7] LadaGAN
@@ IV-H Generated images and interpolation
In addition to the FID-based evaluation, we visually inspect sampled images for a qualitative evaluation. In Figure 6, we present curated images generated by the best performing LadaGAN models for CIFAR-10, CelebA, FFHQ and LSUN bedrooms (see Table II), together with associated attention maps at the $8\times 8$, $16\times 16$, and $32\times 32$ stages, which correspond to the three Ladaformer blocks in Figure 4. In general, we observe that LadaGAN models can generate realistic-looking and diverse images for all datasets. In particular, the images generated by the CIFAR-10 model represent different categories, viewpoints, backgrounds, and even variations within some categories; the CelebA and FFHQ models generate face images with different genres, ethnicities, ages, hairstyles, clothing, viewpoints, and backgrounds; and the LSUN bedroom images contain different styles, colors, and decorations. Interestingly, the $8\times 8$ attention maps resemble access to a single token, whereas the $16\times 16$ and $32\times 32$ maps generate a global structure of the image. Note that for CIFAR-10 and LSUN bedroom the $16\times 16$ maps also seem to converge to a single token, similar to the $8\times 8$ maps, but for FFHQ and CelebA the global structure appears to be preserved. Moreover, as shown in Figure 7, LadaGAN models generate realistic images and smooth transitions of linearly interpolated latent vectors.
@@ V Conclusion
In this paper, we presented LadaGAN, a novel efficient GAN architecture based on a linear additive-attention block called Ladaformer. This block showed to be more suitable for both the generator and the discriminator than other efficient Transformer blocks, allowing stable GAN training in different scenarios. Our findings indicate that Ladaformer is convolution-compatible and LadaGAN is gradient-stable and highly effective for image generation tasks. Remarkably, LadaGAN outperformed ConvNet and Transformer GANs on multiple benchmark datasets at different resolutions while requiring significantly fewer FLOPs. Moreover, compared with diffusion models and Consistency Training, LadaGAN achieves competitive performance at a fraction of the computational cost.

### v5

@@ I Introduction
In recent years, deep generative models have achieved remarkable results in image generation. In particular, Generative Adversarial Networks (GANs) [1] and Diffusion Models (DMs) [2] have become the state-of-the-art approaches for this task. GANs generate images in a single forward pass by learning to map a latent code to realistic samples, whereas diffusion models iteratively refine noise into images using learned denoising processes. Despite their success, GANs and DMs are often computationally expensive, typically requiring millions (and sometimes billions) of parameters and multiple high-end GPUs to train effectively [3, 4]. Moreover, both paradigms involve extensive training iterations: GANs often require prolonged training, while diffusion models are even more costly due to the need to optimize multi-step denoising trajectories across many iterations. This computational burden poses a barrier to accessibility, reproducibility, and rapid experimentation, especially for researchers or developers without access to large-scale infrastructure.
Additionally, training GANs remains notoriously unstable. A large body of research has explored improved objectives (e.g., the Wasserstein loss [5]) and regularization methods (e.g., spectral normalization [6]) to mitigate divergence and mode collapse. Further, state-of-the-art GANs often require laborious engineering and sophisticated neural modules, as seen in convolution-based models like StyleGAN [7, 8], which are computationally demanding in terms of both FLOPs and parameters. Since self-attention has been shown to effectively learn long-range dependencies [9], different GAN architectures that incorporate Transformers [10] have been proposed. However, self-attention can make GAN training even more unstable [11], and its $O(N^{2})$ complexity results in high computational demands [11, 12].
@@
Our key innovations and contributions are as follows:
  • Ladaformer: linear additive attention for stable adversarial training. We introduce Ladaformer, a Transformer block with linear additive attention that enables efficient long-range modeling while remaining stable under adversarial settings. Unlike standard attention, it avoids mode collapse and gradient instabilities common in GANs. Ladaformer is simple, interpretable, and does not require custom kernels or training tricks.
  • LadaGAN: a lightweight, stable Transformer GAN. LadaGAN is designed to enable training from scratch on a single GPU, with significantly reduced training time and computational cost. By integrating Ladaformer blocks into both the generator and discriminator, the architecture achieves high efficiency, requiring far fewer FLOPs and parameters than diffusion-based models, CT, or conventional GANs.
  • Strong performance with minimal compute. LadaGAN achieves competitive or superior FID scores compared to Transformer GANs, diffusion-based models, and CT on CIFAR-10, CelebA, FFHQ, and LSUN Bedroom—without distillation, transfer learning, or large-scale infrastructure. We further benchmark multiple $O(N)$ attention mechanisms under the same low-resource setting, and find that Ladaformer consistently offers the best trade-off between quality and efficiency.
@@ IV Experiments
Attention D-type Convolutions Residual-MLP Modulation G-FLOPs FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ Prec $\uparrow$ Rec $\uparrow$
@@
Conv
Conv
Conv
Conv
Conv
Conv
Lada
@@
0.60
  • “N/A” indicates that training consistently diverged across multiple runs.
@@ IV-C Ablation studies
Table I shows the FID scores, IS Precision, Recall, and number of FLOPs for all the evaluated attention mechanisms and configurations. As can be observed, the Ladaformer with a convolutional layer and without the residual connection achieves the best evaluation with using a convolutoinal discriminator. In general, adding a convolutional layer has a positive effect on Ladaformer generator, a negative effect for Linformer and practically no effect for Fastformer. This shows not only that the combination of Ladaformer and convolutions does not lead to training instabilities, but also that it can provide noticeable benefits for the quality of the generated images. This is because the locality of the convolutional layer might complement the long-range dependencies of the additive attention map that is propagated by the element-wise operation in equation 3. Note that, as opposed to Lada, Fastformer attention mechanism compresses the representation by computing a second additive attention map for the keys instead of propagating it through an element-wise operation, which seems to prevent the benefits of the convolutional layer.
@@
Finally, the best model employs a Ladaformer-based discriminator. Replacing it with Fastformer (listed as “Fast” under D-type in Table I), while keeping almost the same generator and discriminator architecture, leads to large generator weight norms and consistent training divergence. [^3] All Fastformer experiments were performed using the official implementation available at https://github.com/wuch15/Fastformer
@@ IV-D Lada discriminator
Table II shows the FIDs obtained with different configurations of ConvNet and Ladaformer discriminators on CIFAR-10, CelebA and LSUN Bedroom. In addition, the number of FLOPs for the discriminator and generator, as well as the sizes of the embeddings in the generator corresponding to the three Ladaformer blocks in Figure 4, are also presented. As can be observed, the Ladaformer discriminators consistently outperform convolutional discriminators. However, the difference is small for CIFAR-10 and CelebA and significantly larger for LSUN Bedroom where larger generators are employed. Interestingly, a higher learning rate is required for the convolutional discriminator to match the Ladaformer discriminator FID. We also observe that, in LSUN Bedroom, using LEE leads to the lowest FID, while in CIFAR-10 and CelebA LEE does not make any difference. We hypothesize that this behavior is due to the resolution of the dataset, since high-resolution images require larger models. Note that in the LSUN Bedroom experiment, the Ladaformer blocks of the generator increase their embedding dimension locally from $\{1024,256,64\}$ to $\{1024,512,256\}$, which in turn increases the number of FLOPs.
@@ IV-F Comparison with state-of-the-art models
Resolution CT [29] ADM-IP [4, 42] StyleGAN2 [8] VITGAN [11] LadaGAN
@@ IV-H Generated images and interpolation
In addition to the FID-based evaluation, we visually inspect sampled images for a qualitative evaluation. In Figure 6, we present curated images generated by the best performing LadaGAN models for CIFAR-10, CelebA, FFHQ and LSUN bedrooms (see Table II), together with associated attention maps at the $8\times 8$, $16\times 16$, and $32\times 32$ stages, which correspond to the three Ladaformer blocks in Figure 4. In general, we observe that LadaGAN models can generate realistic-looking and diverse images for all datasets. In particular, the images generated by the CIFAR-10 model represent different categories, viewpoints, backgrounds, and even variations within some categories; the CelebA and FFHQ models generate face images with different genres, ethnicities, ages, hairstyles, clothing, viewpoints, and backgrounds; and the LSUN bedroom images contain different styles, colors, and decorations. Interestingly, the $8\times 8$ attention maps resemble access to a single token, whereas the $16\times 16$ and $32\times 32$ maps generate a global structure of the image. Note that for CIFAR-10 and LSUN bedroom the $16\times 16$ maps also seem to converge to a single token, similar to the $8\times 8$ maps, but for FFHQ and CelebA the global structure appears to be preserved. Moreover, as shown in Figures 8 and 7, LadaGAN models generate realistic images and smooth transitions of linearly interpolated latent vectors.
@@
Figure 7: Latent space interpolations and multi-resolution attention maps from the LadaGAN models on FFQH (a, b) and LSUN Bedroom (c, d).
@@ V Conclusion
In this paper, we presented LadaGAN, a novel efficient GAN architecture based on a linear additive-attention block called Ladaformer. This block showed to be more suitable for both the generator and the discriminator than other efficient Transformer blocks, allowing stable GAN training in different scenarios. Our findings indicate that LadaGAN is gradient-stable and highly effective for image generation tasks. Remarkably, LadaGAN outperformed ConvNet and Transformer GANs on multiple benchmark datasets at different resolutions while requiring significantly fewer FLOPs. Moreover, compared with diffusion models and CT, LadaGAN achieves competitive performance at a fraction of the computational cost.